OpenStudy (anonymous):

The point (3, 1/4) is on the graph of y=f(x), and the slope at each pointy (x,y) on the graph is given by dy/dx=y^2(6-2x). Find d^2y/dx^2 and evaluate it at the point (3, 1/4).

OpenStudy (anonymous):

since: \[\frac{dy}{dx}=y^2(6-2x) \rightarrow \frac{dy}{dx}=6y^2-2y^2x\] \[\frac{d^2y}{dx^2}=12y \frac{dy}{dx}-4yx \frac{dy}{dx}-2y^2\] \[\rightarrow \frac{d^2y}{dx^2}=(12y-4yx)\frac{dy}{dx}-2y^2\] now substitute dy/dx in... \[\rightarrow \frac{d^2y}{dx^2}=(12y-4yx)(6y^2-2y^2x)-2y^2\] now plug in 3 for x and 1/4 for y.... let me know what you get

OpenStudy (anonymous):

you should get -1/8